# Zeroth-order log-concave sampling

### v1

@@ 1.1.1 Uniform sampling from a warm start
Kook, Vempala, and Zhang [KVZ24] adapted this idea to convex-body sampling by introducing the $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ algorithm. They extended the contraction result to constrained distributions and developed a new implementation for sampling from $\pi^{X|Y=y}\propto\gamma_{h}(\cdot-y)|_{\mathcal{K}}$. They introduced a threshold parameter $\tau$: if the number of retries in the rejection sampling with proposal $\mathcal{N}(y,hI_{d})$ exceeds $\tau$, then $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ halts and declares failure. To control this failure probability $\eta\in(0,1)$, they carried out involved analysis to carefully set the parameters $h,\eta,\tau$ so that with high probability, the backward step can be completed using $\widetilde{O}(M_{\infty}\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}})$ queries. Consequently, they established $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}\Lambda\operatorname{polylog}\nicefrac{{D}}{{\eta\varepsilon}})$ for $\Lambda:=\lVert\operatorname{cov}\pi\rVert$, generalizing $\mathsf{Ball\ walk}$’s output guarantee. Later, Kook and Zhang [KZ25] established a balanced guarantee of $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ using ($\mathsf{LSI}$), with query complexity $\widetilde{O}(M_{\infty}d^{2}D^{2}\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$.
@@
Nevertheless, algorithmic tricks (e.g., the failure probability) still remain in $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$, and the analysis becomes unwieldy due to intricate interdependencies among the parameters: $h,\eta,\tau$, and the number $N$ of iterations. Moreover, the required initial warmness remains stringent as $\mathcal{R}_{\infty}$. The warmness requirement was later relaxed to $\mathcal{R}_{c}$ with $c=2+\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}}=\widetilde{O}(1)$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{2}$ (i.e., $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ under ($\mathsf{PI}$)), with query complexity $\widetilde{O}(M_{c}d^{2}\Lambda\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ [KV25a]. While this result gave hope for balancing the orders of warmness and final guarantee, the complex parameter dependencies continued to be a major obstacle.
Algorithm Type Complexity Reference
@@
$\mathsf{Ball\ walk}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\operatorname{poly}(d,D,\frac{M_{\infty}}{\varepsilon})$Rightarrow $\frac{M_{\infty}^{2}d^{2}\Lambda}{\varepsilon^{2}}$ [LS93]
$\mathsf{Speedy\ walk}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $M_{\infty}d^{2}D^{2}\log\frac{1}{\varepsilon}$Rightarrow $M_{\infty}d^{2}\Lambda\log\frac{1}{\varepsilon}$ [KLS97]
$\mathsf{Hit\text{-}and\text{-}Run}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\frac{M_{\infty}^{2}d^{2}D^{2}}{\varepsilon^{2}}$ [Lov99]
$\mathcal{R}_{2}\to\mathcal{R}_{2}$ $d^{2}D^{2}\log\frac{M_{2}}{\varepsilon}$ [LV06b]
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ $\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ $qM_{\infty}d^{2}\Lambda\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KVZ24]
$\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ $M_{\infty}d^{2}D^{2}\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KZ25]#
$\mathcal{R}_{\widetilde{O}(1)}\to\mathcal{R}_{2}$ $M_{\widetilde{O}(1)}d^{2}\Lambda\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KV25a]
$\mathsf{PS}_{\textup{unif}}$ $\mathcal{R}_{q}\to\mathcal{R}_{q}$ $qM_{q}^{q/(q-1)}d^{2}\Lambda\log\frac{1}{\varepsilon}$ Theorem 1.1
Table 1: Query complexities of uniform samplers from a convex body. The right arrow ($\Rightarrow$) indicates subsequent improvements due to the improved Poincaré constant of log-concave probability measures. $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ requires the additional parameter $\eta$ referring to a failure probability of the algorithm. The $\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ result in # can be readily extended to $\mathsf{PS}_{\textup{unif}}$ as well, with guarantee $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ of complexity $M_{\infty}d^{2}D^{2}\log\frac{1}{\varepsilon}$ (i.e., removal of $\eta$).
Proximal sampler with restart for balanced guarantees (§2).
Our proposed sampler, $\mathsf{PS}_{\textup{unif}}$—essentially $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ with failure replaced by restart—resolves this issue: starting with $x_{0}\sim\pi_{0}^{X}$, step size $h$, and threshold $\tau$,
  • [Forward] $y_{k+1}\sim\pi^{Y|X=x_{k}}=\gamma_{h}(\cdot-x_{k})$.
[…]
This algorithm allows $h$ and $\tau$ to depend solely on the dimension and initial warmness, leading to the same Rényi orders of initial warmness and final guarantee while leveraging the state-of-the-art bound on Poincaré constants of log-concave probability measures (i.e., $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$)).
Theorem 1.1 (Uniform sampling from a warm start).
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ with $B_{1}\subset\mathcal{K}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$, given access to the membership oracle for $\mathcal{K}$. Let $\pi_{k}$ denote the law of the $k$-th iterate of $\mathsf{PS}_{\textup{unif}}$, initialized from $\pi_{0}$, with step size $h\asymp(d^{2}\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi))^{-1}$, and threshold $\tau\asymp\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}$ for $q\geq 2$. Then, for any $\varepsilon>0$, we can ensure $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$ after $N\asymp qd^{2}\Lambda\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\log d\log\frac{1}{\varepsilon}$ iterations, using
$2\tau N=\widetilde{O}\bigl{(}qd^{2}\Lambda\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}\log\frac{1}{\varepsilon}\bigr{)}\qquad\text{membership queries in expectation}\,.$
Under the warmness notation, the query complexity reads $\widetilde{O}(qd^{2}\Lambda M_{q}^{\frac{q}{q-1}}\log\frac{1}{\varepsilon})$.
1.1.2 Order-preserving warm-start generation via annealing
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ exhibits polynomial dependence on $M_{\widetilde{O}(1)}$, and a “cold start” (e.g., the uniform distribution over a small ball) yields $M_{\widetilde{O}(1)}\lesssim\exp(d,D)$. Hence, warm-start generation is necessary to reduce the complexity overhead incurred by the cold start, so warm-start generating algorithms have been developed alongside sampling algorithms.
Warm-start generation in the well-conditioned setting.
The necessity of a warmness is not limited to this general setting; its importance also arises in the well-conditioned setting (i.e., $-\log\pi$ is $\alpha$-strongly convex and $\beta$-smooth). In fact, $\mathsf{MALA}$ from an $O(1)$-warm start in $\mathcal{R}_{2}$ can take a step size $h\asymp(\beta d^{1/2})^{-1}$, achieving the best first-order complexity of $\widetilde{O}(\kappa d^{1/2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$) in the well-conditioned setting. In contrast, from a feasible start, it can only take a smaller step size $h\asymp(\beta d)^{-1}$, resulting in $\widetilde{O}(\kappa d\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ complexity, where $\kappa:=\beta/\alpha$ denote the condition number of $-\log\pi$. Recently, [AC24] showed that combining the Underdamped Langevin Algorithm with $\mathsf{PS}$ can generate an $O(1)$-warm start in $\mathcal{R}_{2}$-divergence using $\widetilde{O}(\kappa d^{1/2})$ first-order queries, thus incurring no overhead for $\mathsf{MALA}$ as a result.
Simulated annealing.
In our general setup, a classical strategy for warm-start generation is an annealing scheme, which involves a sequence of probability distributions $\{\mu_{i}\}_{i\leq m}$ that interpolate between a tractable initial distribution $\mu_{0}$ and the target distribution $\pi_{\mathcal{K}}$. The key idea is to ensure that each pair of consecutive distributions $\mu_{i}$ and $\mu_{i+1}$ are $O(1)$-close in some appropriate metric (e.g., $\mathcal{R}_{\infty}$-divergence), thereby enabling efficient transitions by a sampler along the sequence.
In the context of convex-body sampling, several annealing schemes have been proposed: (i) uniform distributions over truncated convex bodies intersected with balls of increasing radii [DFK91, LS90, LS93, KLS97], (ii) exponential distributions of the form $e^{-cx}$ with decreasing $c$ over a convex set, derived from the so-called “pencil construction” [LV06c], and (iii) Gaussian distributions with increasing variance, truncated to the body [CV15, CV18, KZ25, KV25a].
Among these, all approaches except [KZ25] provide a warm start $\mu$ satisfying $\lVert\mu-\pi_{\mathcal{K}}\rVert_{\mathsf{TV}}\leq\varepsilon$, while [KZ25] guarantees $\mathcal{R}_{\infty}$-warmness with complexity $\widetilde{O}(d^{2}R^{2})$ for $R^{2}=\mathbb{E}_{\pi_{\mathcal{K}}}[|\cdot|^{2}]$. On the other hand, $\mathsf{TV}$-warmness can be achieved with better complexity $\widetilde{O}(d^{2}R^{3/2}\Lambda^{1/4}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ [KV25a]. Unfortunately, no known result currently interpolates between warmness results in $\mathsf{TV}$ and $\mathcal{R}_{\infty}$ (i.e., either very weak or strong warmness, without known intermediate warmness).
Toward genuine warmness.
First of all, why do we even care about $\mathcal{R}_{q}$-warmness? Given a theoretical gap in known complexities (for generating $\mathcal{R}_{q}$-warmness) between the two settings discussed above, bridging this gap is theoretically intriguing in its own right. In our general setup, it is not even clear whether generating $\mathcal{R}_{q}$-warmness without significant overhead is feasible, let alone whether one can tighten the complexities of warm-start generation and sampling from a warm start.
Moreover, $\mathsf{TV}$-warmness serves only as a “fake” warm-start for downstream sampling in some sense, since subsequent sampling guarantees using this warmness are collapsed down to $\mathsf{TV}$-distance due to the necessity of applying triangle inequalities—an issue referred to as “$\mathsf{TV}$-collapse” by [KZ25]. In practice, this limitation leads to cumulative sampling errors, resulting in somewhat cumbersome post-processing steps, such as replacing the original target accuracy $\varepsilon$ by a reduced accuracy $\varepsilon/m$, where $m$ is the number of times the triangle inequality is applied. Lastly, when combined with $\mathsf{PS}_{\textup{unif}}$ that achieves balanced guarantees, a $\mathcal{R}_{q}$-warmness result would finally yield clean guarantees for uniform sampling in terms of any $\mathcal{R}_{q}$-divergences.
@@ 1.1.2 Order-preserving warm-start generation via annealing
A main difficulty in obtaining $\mathcal{R}_{q}$-warmness through annealing is that we do not yet know how to relay $\mathcal{R}_{q}$-warmness without compromise, leading nearly all previous work to encounter this $\mathsf{TV}$-collapse problem. To illustrate this, consider annealing distributions $\mu_{0},\mu_{1},\mu_{2}$, where a sampler transitions from $\mu_{0}$ toward $\mu_{1}$. MCMC samplers are inherently approximate, producing a measure $\bar{\mu}_{1}$ close to $\mu_{1}$, quantified via $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})\leq\varepsilon$ for some $q>1$. In the next phase, the sampler moves toward $\mu_{2}$ starting from the approximate measure $\bar{\mu}_{1}$ instead of the true $\mu_{1}$. Ideally, the complexity for this phase would directly depend on $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})$. Unfortunately, the lack of an exact triangle inequality for $\mathcal{R}_{q}$-divergence prevents us from ensuring $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})=O(1)$, even if $\mathcal{R}_{q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$.
@@
Boosting a Rényi-order.
The above discussion highlights the fundamental bottleneck arising from the absence of a triangle inequality for Rényi divergences. One promising direction to overcome this is to leverage a weak triangle inequality (1.3), namely, $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})\lesssim\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})+\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})$. As [KV25a] has already designed annealing schemes that can ensure $\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$, one might consider starting with a $\mathcal{R}_{2q-1}$-guarantee when sampling from $\mu_{1}$. However, this effectively amounts to requiring a $\mathcal{R}_{\Theta(2^{m}q)}$-guarantee for $\mu_{1}$-sampling when there are $m$ many phases. Since $\mathsf{PS}$ has linear dependence on $q$, and prior annealing approaches require $m\gtrsim d^{1/2}$, such a “backtracking” analysis would ultimately incur an $\Omega(2^{\sqrt{d}}q)$-multiplicative overhead in the final complexity.
Given the failure of this naïve approach, a more valid question is whether one can ensure $\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})=O(1)$ given $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})<\varepsilon$. Since $2q-1>q$ for $q>1$, reversing the monotonic property of $\mathcal{R}_{q}$-divergence at some overhead will be a main challenge. This reflection naturally introduces an algorithmic question: can we boost Rényi-orders algorithmically (i.e., can we transform $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)=O(1)$ into a stronger bound $\mathcal{R}_{2q-1}(\mu^{\prime}\mathbin{\|}\pi)=O(1)$)? Solving this question would enhance our theoretical toolkit, providing the final piece toward comprehensive convex-body sampling and potentially influencing applications in areas such as differential privacy.

### v2

@@ 1.1.1 Uniform sampling from a warm start
Convex body sampling does not satisfy bounded smoothness. Kook, Vempala, and Zhang [KVZ24] adapted these ideas to convex body sampling by introducing the $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ algorithm. They extended the contraction result to constrained distributions and developed a new efficient implementation for sampling from $\pi^{X|Y=y}\propto\gamma_{h}(\cdot-y)|_{\mathcal{K}}$. They introduced a threshold parameter $\tau$: if the number of attempts in the rejection sampling subroutine with proposal $\mathcal{N}(y,hI_{d})$ exceeds $\tau$, then $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ halts and declares failure. To control this failure probability $\eta\in(0,1)$, they carefully set the parameters $h,\eta,\tau$ so that with high probability, the backward step can be completed using $\widetilde{O}(M_{\infty}\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}})$ queries. Consequently, they established $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}\Lambda\operatorname{polylog}\nicefrac{{D}}{{\eta\varepsilon}})$ for $\Lambda:=\lVert\operatorname{cov}\pi\rVert$, generalizing $\mathsf{Ball\ walk}$’s output guarantee. Later, Kook and Zhang [KZ25] established a balanced guarantee of $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ using ($\mathsf{LSI}$), with query complexity $\widetilde{O}(M_{\infty}d^{2}D^{2}\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$.
@@
Nevertheless, some algorithmic tricks (e.g., the failure probability) still remain in $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$. Moreover, the required initial warmness remains the stringent $\mathcal{R}_{\infty}$. Kook and Vempala [KV25a] relaxed the warmness to $\mathcal{R}_{c}$ with $c=2+\operatorname{polylog}\nicefrac{{dD}}{{\eta\varepsilon}}=\widetilde{O}(1)$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{2}$ (i.e., $\mathsf{\mathsf{In\text{-}and\text{-}Out}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ under ($\mathsf{PI}$)), with query complexity $\widetilde{O}(M_{c}d^{2}\Lambda\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ . This result raised the question and gave hope for fully balancing the warmness and output guarantee.
Proximal sampler with restart for balanced guarantees (§2).
We propose the sampler, $\mathsf{PS}_{\textup{unif}}$—essentially $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ with failure replaced by restart: starting with $x_{0}\sim\pi_{0}^{X}$, step size $h$, and threshold $\tau$,
  • [Forward] $y_{k+1}\sim\pi^{Y|X=x_{k}}=\gamma_{h}(\cdot-x_{k})$.
  • [Backward] $x_{k+1}\sim\pi^{X|Y=y_{k+1}}\propto\gamma_{h}(\cdot-y_{k+1})\,\mathds{1}_{\mathcal{K}}(\cdot)$. To this end, draw $x_{k+1}\sim\mathcal{N}(y_{k+1},hI_{d})$ repeatedly until $x_{k+1}\in\mathcal{K}$. If this rejection loop exceeds $\tau$ trials, declare failure and restart scratch from a new $x_{0}\sim\pi_{0}^{X}$.
This algorithm always succeeds and has a complexity guarantee from an $O(1)$-warm start, where the Rényi divergence orders of initial warmness and output distribution guarantee are the same (i.e., $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$)). This is our first main result. Hereafter, we use $B_{1}$, $B_{2}$, or $B_{2}^{d}$ to denote a $\ell_{2}$-unit ball in $\mathbb{R}^{d}$.
Theorem 1.1 (Uniform sampling from a warm start).
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ given by a membership oracle with $B_{1}\subset\mathcal{K}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$. Given $\varepsilon\in(0,1/100)$, an initial distribution $\pi_{0}$, and $q\geq 2\vee\widetilde{\Omega}(\log(d^{2}\Lambda\log\frac{1}{\varepsilon}))$ such that $M_{q}=\lVert\nicefrac{{\mathrm{d}\pi_{0}}}{{\mathrm{d}\pi}}\rVert_{L^{q}(\pi)}\leq 10$, consider the algorithm that runs $\mathsf{PS}_{\textup{unif}}$ initialized at $\pi_{0}$ and restarts if $\mathsf{PS}_{\textup{unif}}$ fails before completing $N=\widetilde{O}(qd^{2}\Lambda\log^{2}\frac{1}{\varepsilon})$ iterations. There exist choices of $h$ and $\tau$ such that this algorithm restarts with probability at most $1/50$, and its output $Z$ satisfies $\mathcal{R}_{q}(\operatorname{law}Z\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(qd^{2}\Lambda\log^{6}\frac{1}{\varepsilon})$ total membership queries in expectation.
1.1.2 Improved warm-start generation via annealing
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ exhibits polynomial dependence on $M_{\widetilde{O}(1)}$, and a “cold” start (e.g., the uniform distribution over a small ball) yields $M_{\widetilde{O}(1)}\lesssim\exp(d,D)$. Hence, efficient warm-start generation is necessary to reduce the complexity overhead incurred by a cold start, so warm-start generation algorithms have been developed alongside sampling algorithms.
Warm-start generation in the well-conditioned setting.
The necessity of warmness is not limited to this general setting; its importance also arises in the well-conditioned setting (i.e., $\log\pi$ is $\alpha$-strongly convex and $\beta$-smooth). In fact, $\mathsf{MALA}$ from an $O(1)$-warm start in $\mathcal{R}_{2}$ can take a step size $h\asymp(\beta d^{1/2})^{-1}$, achieving the first-order complexity of $\widetilde{O}(\kappa d^{1/2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$) in the well-conditioned setting. In contrast, from a feasible start, it can only take a smaller step size $h\asymp(\beta d)^{-1}$, resulting in $\widetilde{O}(\kappa d\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ complexity, where $\kappa:=\beta/\alpha$ denote the condition number of $\log\pi$. Recently, [AC24] showed that combining the Underdamped Langevin Algorithm with $\mathsf{PS}$ can generate an $O(1)$-warm start in $\mathcal{R}_{2}$-divergence using $\widetilde{O}(\kappa d^{1/2})$ first-order queries, thus incurring no overhead for $\mathsf{MALA}$ as a result.
Simulated annealing.
In our general setup, a classical strategy for warm-start generation is an annealing scheme, which involves a sequence of probability distributions $\{\mu_{i}\}_{i\leq m}$ to interpolate between an easy-to-sample initial distribution $\mu_{0}$ and the target distribution $\pi_{\mathcal{K}}$. The key idea is to ensure that each pair of consecutive distributions $\mu_{i}$ and $\mu_{i+1}$ is $O(1)$-close in some appropriate metric (e.g., $\mathcal{R}_{\infty}$-divergence), thereby enabling efficient transitions by a sampler along the sequence.
In the context of convex-body sampling, several annealing schemes have been proposed: (i) uniform distributions over truncated convex bodies intersected with balls of increasing radii [DFK91, LS90, LS93, KLS97], (ii) exponential distributions of the form $e^{-cx}$ with decreasing $c$ over a convex set, derived from the so-called “pencil” construction [LV06c], and (iii) Gaussian distributions with increasing variance, truncated to the body [CV15, CV18, KZ25, KV25a].
Among these, all approaches except [KZ25] provide a warm start $\mu$ satisfying $\lVert\mu-\pi_{\mathcal{K}}\rVert_{\mathsf{TV}}\leq\varepsilon$, while [KZ25] guarantees $\mathcal{R}_{\infty}$-warmness with complexity $\widetilde{O}(d^{2}R^{2})$ for $R^{2}=\mathbb{E}_{\pi_{\mathcal{K}}}[|\cdot|^{2}]$. [KV25a] show that $\mathsf{TV}$-warmness can be achieved with the better complexity of $\widetilde{O}(d^{2}R^{3/2}\Lambda^{1/4}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}})$ . Currently, these are the only known guarantees, and there is no bound for intermediate warmness results, between $\mathsf{TV}$ and $\mathcal{R}_{\infty}$ (i.e., either very weak or strong warmness).
Towards genuine warmness.
Why do we even care about $\mathcal{R}_{q}$-warmness? Given the gap in known complexities between the two extreme settings discussed above, bridging this gap is theoretically intriguing in its own right. In our general setup, it is not clear whether generating $\mathcal{R}_{q}$-warmness without significant overhead is feasible, let alone whether one can tighten the complexities of warm-start generation and sampling from a warm start.
Moreover, $\mathsf{TV}$-warmness arguably serves only as a “pseudo” warm-start for downstream sampling, since subsequent sampling guarantees using this warmness are only in $\mathsf{TV}$-distance due to the necessity of applying triangle inequalities. In practice, this limitation leads to accumulation of sampling errors, resulting in somewhat cumbersome post-processing, such as replacing the original target accuracy $\varepsilon$ by a reduced accuracy $\varepsilon/m$, where $m$ is the number of times the triangle inequality is applied. While this does not significantly affect downstream guarantees for applications (since $\varepsilon^{-1}$ is inside a logarithm), it is a bit unsatisfactory and makes the analysis more complicated.
One might instead generate a $\mathcal{R}_{\infty}$-warm start and then apply the known $\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ guarantee (see Figure 1.1). However, generating $\mathcal{R}_{\infty}$-warmness incurs a higher query complexity than $\mathsf{TV}$-warmness, making it an unnecessarily expensive route to an $\mathcal{R}_{q}$-guarantee. Lastly, when combined with $\mathsf{PS}_{\textup{unif}}$, which achieves balanced guarantees, a $\mathcal{R}_{q}$-warmness result would yield clean end-to-end guarantees for uniform sampling in any desired $\mathcal{R}_{q}$-divergence.
@@ 1.1.2 Improved warm-start generation via annealing
Figure 1.1: Zeroth-order query complexities for logconcave sampling. Arrows indicate the sampling complexity from an $O(1)$-warm start. Complexities listed next to each left node indicate warmness-generation costs. The dagger sign ($\dagger$) indicates results for uniform sampling over convex bodies (membership oracle). Results without $\dagger$ hold for general logconcave distributions (evaluation oracle). Here, $\bar{R}:=R\vee 1$, $\bar{\lambda}:=\lambda\vee 1$, $\lambda:=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}:=\mathbb{E}_{\pi}[|\cdot|^{2}]$ (so $\lambda\leq R^{2}$). Note that $\mathsf{TV}$-warmness can be treated as $\mathcal{R}_{\infty}$-warmness at the cost of collapsing downstream guarantees to $\mathsf{TV}$-distance.
@@
A central difficulty in obtaining $\mathcal{R}_{q}$-warmness through annealing is that we do not yet know how to relay $\mathcal{R}_{q}$-warmness without compromise, leading nearly all previous work to encounter this $\mathsf{TV}$-collapse problem. To illustrate this, consider the annealing distributions $\mu_{0},\mu_{1},\mu_{2}$, where a sampler transitions from $\mu_{0}$ toward $\mu_{1}$. Markov-chain samplers for convex bodies are inherently approximate, producing a measure $\bar{\mu}_{1}$ close to but not exactly $\mu_{1}$, quantified via $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})\leq\varepsilon$ for some $q>1$. In the next phase, the sampler moves toward $\mu_{2}$ starting from the approximate measure $\bar{\mu}_{1}$, not $\mu_{1}$. Ideally, the complexity for this phase would directly depend on $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})$. Unfortunately, the lack of a triangle inequality for $\mathcal{R}_{q}$-divergence prevents us from ensuring $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})=O(1)$, even if $\mathcal{R}_{q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$.
@@
Boosting the Rényi divergence order.
The above discussion highlights the fundamental bottleneck arising from the absence of a triangle inequality for Rényi divergences. One promising direction to overcome this is to leverage a weak triangle inequality (1.3), namely, $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{2})\lesssim\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})+\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})$. As [KV25a] has already designed annealing schemes that can ensure $\mathcal{R}_{2q}(\mu_{1}\mathbin{\|}\mu_{2})=O(1)$, one might consider starting with a $\mathcal{R}_{2q-1}$-guarantee when sampling from $\mu_{1}$. However, this effectively amounts to requiring a $\mathcal{R}_{\Theta(2^{m}q)}$-guarantee for $\mu_{1}$-sampling when there are $m$ phases. Since $\mathsf{PS}$ has linear dependence on $q$, and prior annealing approaches require $m\gtrsim d^{1/2}$, such a “backtracking” analysis would potentially incur an $\Omega(2^{\sqrt{d}}q)$-multiplicative overhead in the final complexity.
Given the failure of this naïve approach, a pertinent question is whether one can ensure $\mathcal{R}_{2q-1}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})=O(1)$ given $\mathcal{R}_{q}(\bar{\mu}_{1}\mathbin{\|}\mu_{1})<\varepsilon$. Since 2q1>q2q-1>q for $q>1$, this seemingly goes in the wrong direction — opposite the monotonic property of $\mathcal{R}_{q}$-divergence. This suggests an algorithmic question: can we boost the Rényi-divergence order algorithmically (i.e., can we transform $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)=O(1)$ into a stronger bound $\mathcal{R}_{2q-1}(\mu^{\prime}\mathbin{\|}\pi)=O(1)$)? Solving this question would enhance our theoretical toolkit, providing the final missing piece towards a comprehensive approach for convex body sampling and its applications.